# Quantifying and Bridging the Fidelity Gap: A Decisive-Feature Approach to Comparing Synthetic and Real Imagery

### v1

@@ Abstract
Virtual testing using synthetic data has become a cornerstone of autonomous vehicle (AV) safety assurance. Despite progress in improving visual realism through advanced simulators and generative AI, recent studies reveal that pixel-level fidelity alone does not ensure reliable transfer from simulation to the real world. What truly matters is whether the system-under-test (SUT) bases its decisions on the same causal evidence in both real and simulated environments—not just whether images “look real” to humans.
This paper addresses the lack of such a behavior-grounded fidelity measure by introducing Decisive Feature Fidelity (DFF), a new SUT-specific metric that extends the existing fidelity spectrum to capture mechanism parity—the agreement in causal evidence underlying the SUT’s decisions across domains.
DFF leverages explainable-AI (XAI) methods to identify and compare the decisive features driving the SUT’s outputs for matched real–synthetic pairs. We further propose practical estimators based on counterfactual explanations, along with a DFF-guided calibration scheme to enhance simulator fidelity.
Experiments on 2126 matched KITTI–VirtualKITTI2 pairs demonstrate that DFF reveals discrepancies overlooked by conventional output-value fidelity. Furthermore, results show that DFF-guided calibration improves decisive-feature and input-level fidelity without sacrificing output value fidelity across diverse SUTs.
@@ I Introduction
Scenario-based virtual testing with synthetic data is a standard for AV safety evaluation [iso_scenario_2022, bsi_scenario_2023], enabling the recreation of complex driving situations.
Substantial resources are being directed toward the development of high-fidelity simulators [dosovitskiyCARLAOpenUrban2017] and generative-AI techniques [rombachHighResolutionImageSynthesis2022, zhangAddingConditionalControl2023] to enhance realism to better reflect actual driving conditions. Most existing efforts have concentrated on improving “pixel-level realism”, that is, aligning synthetic image data with real image data in a way that is agnostic to the system-under-test (SUT). For brevity we focus on image-based data and refer to pixels when discussing input-level fidelity, but the proposed approach extends to other modalities such as LiDAR by replacing pixel correspondence with point-cloud correspondence.
However, recent studies suggest that pixel-level, SUT-agnostic fidelity does not assure reliable transfer of AV assessment conclusions between the virtual and real domains. What truly matters is whether the AV grounds its behaviors or decisions on the same underlying principles in both real and synthetic domains[chengInstanceLevelSafetyAwareFidelity2024, stoccoMindGapStudy2023, johnsonLiteratureReviewSimulation2023, zhao_statistical_2025]. This emerging consensus has shifted attention from pixel-level realism of the data towards behavioral consistency of the SUT as a more meaningful indicator of simulation validity [Riedmaier2020Survey, Riedmaier2022Taxonomy, Tang2023ADSTesting, Omeiza2022XAI].
@@
Figure 1: Top: three levels of the fidelity spectrum (IV, LF, OV) from [chengInstanceLevelSafetyAwareFidelity2024]. Bottom: our extension that adds decisive-feature fidelity (DFF) and a calibrator to adjust the generator. Safety-aware fidelity (SA) is defined in [chengInstanceLevelSafetyAwareFidelity2024] but is not explicitly visualized here.
The top part of Fig. 1 recalls the input/latent/output/safety-aware fidelity spectrum of [chengInstanceLevelSafetyAwareFidelity2024]. The bottom part illustrates our extension: an XAI explainer extracts decisive features from the SUT for each paired real and synthetic input, the distance between decisive maps yields a DFF estimate, and this estimate is fed back to a calibrator that adjusts generator knobs to close mechanism gaps. The previous work [chengInstanceLevelSafetyAwareFidelity2024] systematically categorized synthetic-data fidelity into four levels, as shown in Fig. 1:
@@
input-level (IV) fidelity, latent-feature (LF) fidelity, output-value (OV) fidelity, and safety-aware (SA) fidelity. As the names suggest, these correspond respectively to the fidelity measured in the SUT’s input space (e.g., the perception neural network’s raw inputs), latent feature space, output vector space, and safety-critical output space. While this spectrum of fidelity spans the entire processing pipeline of a neural network, it remains too coarse to capture why the SUT exhibits the same or different behaviors when presented with paired real and synthetic inputs representing the same scenario. For example, an AV perception module may output the same steering command, say a 20-degree left turn, for both the real and synthetic inputs, implying high OV-fidelity. Yet the underlying reasons could differ: the real image might trigger the action due to a detected dog ahead, whereas in the synthetic version of the same scenario, the same decision arises from detecting a right-turn sign next to the dog (and the dog is misdetected, which is concerning). In this case, the system produces the same decision for different reasons, i.e., based on different causal evidence. Notably, the LF-fidelity defined in [chengInstanceLevelSafetyAwareFidelity2024] cannot fully capture such discrepancies, as it simply compares all extracted features without distinguishing those that actually influence the SUT’s decision. This reveals a gap: existing metrics tell us what the SUT outputs, but not why. We need a metric that checks whether the SUT uses the same visual evidence (e.g., road edges, lane markings) in both domains. This gap directly motivates DFF. Throughout this paper, we evaluate and calibrate synthetic data for testing a fixed SUT; we do not train or adapt the SUT itself.
To bridge this gap, we propose an explainable-AI (XAI)-based fidelity metric that quantifies the differences in the decisive features influencing the SUT’s outputs for a given pair of real and synthetic data. Specifically, we first formalize a new SUT-specific fidelity metric, termed Decisive Feature Fidelity (DFF), which extends the fidelity spectrum defined in [chengInstanceLevelSafetyAwareFidelity2024].
Unlike the LF-fidelity in [chengInstanceLevelSafetyAwareFidelity2024], which compares all extracted features, DFF uses recent XAI methods to estimate and compare only the (subset of) features that truly influence the SUT’s output.
We then design practical estimators based on counterfactual explanations to quantify DFF and use these estimates as feedback signals to recalibrate the synthetic data generator, thereby improving fidelity.
In our experiments we instantiate $G$ with a diffusion-based generator, but the proposed calibration objective is general and applies equally to reconfigurable simulators such as CARLA or PreScan.
To validate our approach, we conduct experiments on 2126 matched KITTI–VirtualKITTI2 pairs across three SUTs (steering, YOLOP drivable-area, and YOLOP lane-line), demonstrating the utility of DFF for both assessment and calibration.
@@ II Related Work
Virtual testing offers a scalable and risk-free way to assess AV safety across diverse scenarios. However, the fidelity of synthetic data used in such testing remains a critical and actively studied issue due to the persistent sim-to-real gap [10242366, 10588858, dosovitskiyCARLAOpenUrban2017, geigerAreWeReady2012]. While fidelity can be defined and classified in multiple ways, an important distinction lies between SUT-agnostic and SUT-specific measures. This perspective sets the foundation for the subsequent review of related works.
@@ II-A SUT-Agnostic Approaches to Fidelity
Domain randomization applies random perturbations (lighting, textures, object placement) during training to encourage models to ignore irrelevant variation; however, it can inadvertently remove task-critical structure that the SUT needs [tobin2017domain]. Game-engine and physically-based rendering (PBR) synthesis can generate large-scale, diverse scenes but does not guarantee that the SUT will process synthetic and real images through equivalent internal pathways [richter2016playing, cabonVirtualKITTI22020]. Image-translation methods such as SimGAN and CycleGAN learn mappings between synthetic and real domains to reduce visual discrepancy; yet without explicit semantic constraints, they may alter safety-relevant content (e.g., removing a pedestrian or shifting lane boundaries) [shrivastava2017simgan, zhu2017cyclegan]. Diffusion models with spatial control (e.g., ControlNet) condition generation on semantic layouts or edge maps, improving structural fidelity to scene geometry; however, they do not verify whether the SUT’s internal reasoning transfers [rombachHighResolutionImageSynthesis2022, zhangAddingConditionalControl2023]. Metric analyses further caution that popular perceptual scores (LPIPS, FID) correlate poorly with downstream task performance [Zhang2018LPIPS]. While valuable, all these methods are SUT-agnostic: they improve visual similarity without checking whether the SUT relies on the same causal evidence. This limitation motivates SUT-specific approaches.
@@ II-B SUT-Specific Fidelity Metrics
Recognizing the limitations of agnostic metrics, some work has moved towards SUT-specific fidelity. Studies like [haq2021can, codevilla2018offline] measured fidelity by comparing the SUT’s performance on real and synthetic data, which essentially corresponds to the OV-fidelity defined in [chengInstanceLevelSafetyAwareFidelity2024]. Beyond OV-fidelity, [chengInstanceLevelSafetyAwareFidelity2024] further introduced a spectrum of fidelity metrics spanning the input, hidden, and output layers of the SUT, assuming a neural network architecture. Moreover, it formalizes instance-level safety-aware fidelity (SA) and shows that calibrating generators to safety outcomes (i.e., those outputs that have actual safety effects on the whole AV system) aligns better with the purpose of AV testing. However, none of these metrics capture mechanism parity—whether the SUT made its decision for the same underlying reasons, not just whether it produced the same final output. This gap motivates the need for a decisive-feature-level metric.
Existing Fidelity Metric Blind Spots
As summarized in Table I, these established metrics have known blind spots. IV fidelity fails when the SUT is locally unrobust: small input perturbations invisible to pixel metrics can cause large output changes [Szegedy2014Intriguing, chengInstanceLevelSafetyAwareFidelity2024], so high IV-fidelity does not guarantee high OV-fidelity. OV fidelity can mask mechanism shifts: the SUT may produce identical outputs from different causal evidence (spurious correlation). LF fidelity, which compares broad internal activations, conflates decisive and non-decisive features [Adebayo2018SanityChecks]; a synthetic image may activate similar neurons overall yet rely on entirely different evidence for the final decision. These gaps motivate a metric that specifically targets the decisive feature subspace—the subset of features that actually influence the SUT’s output.
@@ III Foundation
We denote the SUT by a function $F:\mathbb{R}^{d_{0}}\to\mathbb{R}^{d_{L}}$ mapping $d_{0}$-dimensional inputs to $d_{L}$-dimensional outputs, and the XAI explainer by $\operatorname{H}$. We let $x_{r}$ and $x_{s}$ denote real and synthetic inputs, respectively. $\operatorname{D}$ represents the distance metric used to quantify the difference between explanations. Decisive features are evaluated through the maps $\operatorname{H}(F(\cdot))$ and compared using $\operatorname{D}$. (All notations are summarized in Appendix A.)
@@ III-A Scenario Description and Generator
Let $\mathbb{A}$ denote the universe of semantic attributes (e.g., actors and environment factors). A scenario description is $SD=\{(a_{k},v_{k})\}_{k=1}^{K}$ with $a_{k}\in\mathbb{A}$ and values $v_{k}$ (discrete or continuous, encoded in $\mathbb{R}$ as needed). A synthetic generator $G_{\Theta}$ takes $SD$ and produces a sample $x_{s}\!=\!G_{\Theta}(SD)\in\mathbb{R}^{d_{0}}$; $G_{\Theta}$ may be stochastic. Let $\mathrm{RW}\!\bigl(SD\bigr)$ denote the (ideal) set of real samples consistent with $SD$. In practice one only has a collected dataset $X^{cr}$ and an empirical approximation $\widehat{\mathrm{RW}}(SD)=X^{cr}\cap\mathrm{RW}(SD)$.
@@ III-B Established Fidelity Metrics
We introduce the formal definitions of fidelity metrics from [chengInstanceLevelSafetyAwareFidelity2024] that are related to our DFF in this subsection. Let $\operatorname{D_{\mathrm{in}}}$ be a distance on $\mathbb{R}^{d_{0}}$, $\operatorname{D_{\mathrm{out}}}$ a distance on $\mathbb{R}^{d_{L}}$, and let $\varepsilon_{\mathrm{in}},\varepsilon_{\mathrm{out}}>0$.
@@
A synthetic $x_{s}\sim G_{\Theta}(SD)$ is $\langle\operatorname{D_{\mathrm{in}}},\varepsilon_{\mathrm{in}}\rangle$-IV-fidelitous if
@@
Given $F$, $x_{s}\sim G_{\Theta}(SD)$ is $\langle F,\operatorname{D_{\mathrm{out}}},\varepsilon_{\mathrm{out}}\rangle$-OV-fidelitous if
@@
Beyond input/output matching, one can assess whether the SUT processes synthetic inputs through similar internal mechanisms. Given a deep neural network (DNN) $F$ with intermediate layers indexed by $\ell$, and using a distance $\operatorname{D}_{\mathrm{lf}}$ and tolerance $\varepsilon_{\mathrm{lf}}$:
@@
IV Proposed Solution: Decisive-Feature Fidelity (DFF)
Our proposed framework, illustrated in Fig. 1, consists of three main components: (1) Metric Definition: We formalize DFF to quantify mechanism parity alongside IV and OV fidelity. (2) Estimation: An XAI explainer extracts decisive features from both real and synthetic inputs to compute the DFF metric. (3) Calibration: A feedback loop uses the DFF metric to optimize the generator’s parameters $\Theta$, minimizing the mechanism gap. The following subsections detail each component.
@@ IV-A DFF Metric Definition
Note that we exclude LF fidelity ($\varepsilon_{\text{lf}}$) from the top-level acceptance criteria as DFF provides a more specific measure of internal mechanism alignment relevant to the task. The thresholds $\varepsilon_{\text{in}}$, $\varepsilon_{\text{out}}$, and $\varepsilon_{\text{dff}}$ are user-defined hyperparameters. In practice, they are derived empirically from a calibration set (e.g., using the 95th percentile of distances between matched real pairs) or set based on industry safety standards (e.g., OV error $5\%$). We instantiate decisive features via an XAI explainer $\operatorname{H}$ and compare them with a distance $\operatorname{D}$. Formally, $x_{s}$ is $\langle F,\operatorname{H},\operatorname{D},\varepsilon_{\text{dff}}\rangle$-DFF-fidelitous with respect to $x_{r}$ if
@@
$\text{Pass-Rate}=\frac{1}{N}\sum_{i=1}^{N}\mathbb{I}\Bigl(\operatorname{D}\bigl(\operatorname{H}(F(x_{s,i})),\,\operatorname{H}(F(x_{r,i}))\bigr)\leq\varepsilon_{\text{dff}}\Bigr).$
@@
IV enforces pixel proximity; OV enforces output proximity; LF enforces broad feature proximity. DFF enforces decisive feature alignment and can diverge from OV/LF when mechanism shifts occur (e.g., correct output for the wrong reason).

### v3

# Quantifying Fidelity: A Decisive Feature Approach to Comparing Synthetic and Real Imagery

@@ Abstract
Virtual testing using synthetic data has become a cornerstone of autonomous vehicle (AV) safety assurance. Despite progress in improving visual realism through advanced simulators and generative AI, recent studies reveal that pixel-level fidelity alone does not ensure reliable transfer from simulation to the real world. What truly matters is whether the system-under-test (SUT) bases its decisions on consistent decision evidence in both real and simulated environments, not just whether images “look real” to humans. To this end this paper proposes a behavior-grounded fidelity measure by introducing Decisive Feature Fidelity (DFF), a new SUT-specific metric that extends the existing fidelity spectrum to capture mechanism parity, that is, agreement in the model-specific decisive evidence that drives the SUT’s decisions across domains. DFF leverages explainable-AI methods to identify and compare the decisive features driving the SUT’s outputs for matched real–synthetic pairs. We further propose estimators based on counterfactual explanations, along with a DFF-guided calibration scheme to enhance simulator fidelity. Experiments on 2126 matched KITTI–VirtualKITTI2 pairs demonstrate that DFF reveals discrepancies overlooked by conventional output-value fidelity. Furthermore, results show that DFF-guided calibration improves decisive-feature and input-level fidelity without sacrificing output value fidelity across diverse SUTs.
@@ I Introduction
Scenario-based virtual testing with synthetic data is a standard for AV safety evaluation [13, 2], enabling the recreation of complex driving situations. Substantial resources are being directed toward high-fidelity simulators [6] and generative-AI methods [20, 26] to improve realism. Most existing efforts focus on “pixel-level realism”, aligning synthetic images with real images in a way that is agnostic to the system-under-test (SUT). However, recent studies indicate that pixel-level similarity does not ensure reliable transfer of assessment conclusions between virtual and real domains [4, 22, 14, 28]. What matters is whether the SUT bases its decisions on consistent visual evidence across domains. This has shifted attention from appearance-based fidelity toward behavior-grounded validity and SUT-specific fidelity notions [18, 19, 24, 16]. In this paper, we focus on image inputs for brevity, but the framework generalizes to other modalities by replacing pixel correspondence with modality-appropriate correspondence.
@@
Figure 1: Top: three levels of the fidelity spectrum (IV, LF, OV) from [4]. Bottom: our extension that adds decisive-feature fidelity (DFF) and a calibrator to adjust the generator. Safety-aware fidelity (SA) is defined in [4] but is not explicitly visualized here.
@@
Previous work [4] systematically categorized synthetic-data fidelity into four levels, as shown in Fig. 1: input-level (IV) fidelity, latent-feature (LF) fidelity, output-value (OV) fidelity, and safety-aware (SA) fidelity. As the names suggest, these correspond respectively to the fidelity measured in the SUT’s input space (e.g., the perception neural network’s raw inputs), latent feature space, output vector space, and safety-critical output space. While this spectrum of fidelity spans the entire processing pipeline of a neural network, it is too coarse to capture why the SUT exhibits the same or different behaviors when presented with paired real and synthetic inputs representing the same scenario. For example, an AV perception module may output the same steering command, say a 20-degree left turn, for both the real and synthetic inputs, implying high OV-fidelity. Yet the underlying reasons could differ: the real image might trigger the action due to a detected pedestrian ahead, whereas in the synthetic version of the same scenario, the same decision arises from detecting a right-turn sign next to the dog (and the dog is misdetected, which is concerning). In this case, the system produces the same decision for different reasons, i.e., based on different causal evidence. Notably, the LF-fidelity defined in [4] cannot fully capture such discrepancies, as it simply compares all extracted features without distinguishing those that actually influence the SUT’s decision. This reveals a gap: existing metrics tell us what the SUT outputs, but not why. We need a metric that checks whether the SUT uses the same visual evidence (e.g., road edges, lane markings) in both domains. This gap directly motivates DFF. The top part of Fig. 1 recalls the input/latent/output/safety-aware fidelity spectrum of [4]. The bottom part illustrates our extension: an Explainable AI (XAI) explainer extracts decisive features from the SUT for each paired real and synthetic input, the distance between decisive maps yields a Decisive Feature Fidelity (DFF) estimate, and this estimate is feedback to a calibrator that adjusts generator parameters to close mechanism gaps. Throughout this paper, we evaluate and calibrate synthetic data for testing a fixed SUT; we do not train or adapt the SUT itself. To bridge this gap, we propose an explainable-AI (XAI)-based fidelity metric that quantifies the differences in the decisive features influencing the SUT’s outputs for a given pair of real and synthetic data. Specifically, we first formalize a new SUT-specific fidelity metric, termed Decisive Feature Fidelity (DFF), which extends the fidelity spectrum defined in [4]. Unlike the LF-fidelity in [4], which compares all extracted features, DFF uses recent XAI methods to estimate and compare only the (subset of) features that truly influence the SUT’s output. We then design practical estimators based on counterfactual explanations to quantify DFF and use these estimates as feedback signals to recalibrate the synthetic data generator, thereby improving fidelity. In our experiments we use a diffusion-based image generator, but the proposed calibration objective is general and also applies to reconfigurable simulators such as CARLA or Prescan. To validate our apphroach, we conduct experiments on 2126 matched KITTI–VirtualKITTI2 pairs across three SUTs (steering, YOLOP drivable-area, and YOLOP lane-line), showing the utility of DFF for both assessment and calibration. DFF is most appropriate in clone/regression regimes where the simulator aims to reproduce a real scene (or a nahrrowly defined scenario description) and the goal is to ensure that synthetic test conclusions transfer to the corresponding real setting. In counterfactual stress testing regimes, simulation is intentionally modified (e.g., actors moved, hazards injected, weather changed), and strict alignment to a particular real frame is not necessarily desirable. In such cases, DFF is better used as (i) a within-regime consistency check (mechanism stability across controlled perturbations), or (ii) a distributional constraint to keep decisive evidence within an acceptable envelope rather than matching a specific real instance.
@@ II Related Work
Virtual testing offers a scalable and risk-free way to assess AV safety across diverse scenarios. However, the fidelity of synthetic data used in such testing remains a critical and actively studied issue due to the persistent sim-to-real gap [12, 15, 6, 8]. While fidelity can be defined and classified in multiple ways, an important distinction lies between SUT-agnostic and SUT-specific measures. This perspective sets the foundation for the subsequent review of related works.
@@ II-A SUT-Agnostic Approaches to Fidelity
Domain randomization applies random perturbations (lighting, textures, object placement) during training to encourage models to ignore irrelevant variation; however, it can inadvertently remove task-critical structure that the SUT needs [25]. Game-engine and physically-based rendering (PBR) synthesis can generate large-scale, diverse scenes but does not guarantee that the SUT will process synthetic and real images through equivalent internal pathways [17, 3]. Image-translation methods such as SimGAN and CycleGAN learn mappings between synthetic and real domains to reduce visual discrepancy; yet without explicit semantic constraints, they may alter safety-relevant content (e.g., removing a pedestrian or shifting lane boundaries) [21, 30]. Diffusion models with spatial control (e.g., ControlNet) condition generation on semantic layouts or edge maps, improving structural fidelity to scene geometry; however, they do not verify whether the SUT’s internal reasoning transfers [20, 26]. Metric analyses further caution that popular perceptual scores (LPIPS, FID) correlate poorly with downstream task performance [27]. While valuable, all these methods are SUT-agnostic: they improve visual similarity without checking whether the SUT relies on the same causal evidence motivating SUT-specific approaches.
@@ II-B SUT-Specific Fidelity Metrics
Recognizing the limitations of agnostic metrics, some work has moved towards SUT-specific fidelity. Studies like [10, 5] measured fidelity by comparing the SUT’s performance on real and synthetic data, which essentially corresponds to the OV-fidelity defined in [4]. Beyond OV-fidelity, [4] further introduced a spectrum of fidelity metrics spanning the input, hidden, and output layers of the SUT, assuming a neural network architecture. Moreover, it formalizes instance-level safety-aware fidelity (SA) and shows that calibrating generators to safety outcomes (i.e., those outputs that have actual safety effects on the whole AV system) aligns better with the purpose of AV testing. However, none of these metrics capture mechanism parity (whether the SUT made its decision for the same underlying reasons, not just whether it produced the same final output. This necessitates a decisive-feature metric).
As summarized in Table I, these established metrics have known blind spots. IV fidelity fails when the SUT is locally unrobust: small input perturbations invisible to pixel metrics can cause large output changes [23, 4], so high IV-fidelity does not guarantee high OV-fidelity. OV fidelity can mask mechanism shifts: the SUT may produce identical outputs from different causal evidence (spurious correlation). LF fidelity, which compares broad internal activations, conflates decisive and non-decisive features [1]; a synthetic image may activate similar neurons overall yet rely on entirely different evidence for the final decision. These gaps motivate a metric that specifically targets the decisive feature subspace, that is, the subset of features that actually influence the SUT’s output.
@@ III Foundation
We denote the SUT by a function $F:\mathbb{R}^{d_{0}}\to\mathbb{R}^{d_{L}}$ mapping $d_{0}$-dimensional inputs to $d_{L}$-dimensional outputs, and the XAI explainer by $\operatorname{H}$. We let $x_{r}$ and $x_{s}$ denote real and synthetic inputs, respectively. $\operatorname{D}$ represents the distance metric used to quantify the difference between explanations. Decisive features are evaluated through the maps $\operatorname{H}(F(\cdot))$ and compared using $\operatorname{D}$.
@@ III-A Scenario Description and Generator
Let $\mathbb{A}$ denote the universe of semantic attributes (e.g., actors and environment factors). A scenario description is $SD=\{(a_{k},v_{k})\}_{k=1}^{K}$ with attributes $a_{k}\in\mathbb{A}$ and values $v_{k}$ (discrete or continuous, encoded in $\mathbb{R}$ as needed). A synthetic generator $G_{\Theta}$ takes $SD$ and produces a sample $x_{s}\!=\!G_{\Theta}(SD)\in\mathbb{R}^{d_{0}}$; $G_{\Theta}$ may be stochastic. Let $\mathrm{RW}\!\bigl(SD\bigr)$ denote the (ideal) set of real samples consistent with $SD$. In practice one only has a collected dataset $X^{cr}$ and an empirical approximation $\widehat{\mathrm{RW}}(SD)=X^{cr}\cap\mathrm{RW}(SD)$.
@@ III-B Established Fidelity Metrics
We introduce the formal definitions of fidelity metrics from [4] that are related to our DFF in this subsection. Let $\operatorname{D_{\mathrm{in}}}$ and $\operatorname{D_{\mathrm{out}}}$ compute the distance on $\mathbb{R}^{d_{0}}$ and $\mathbb{R}^{d_{L}}$, and let $\varepsilon_{\mathrm{in}},\varepsilon_{\mathrm{out}}>0$.
@@
A synthetic $x_{s}\!=\!G_{\Theta}(SD)$ is $\langle\operatorname{D_{\mathrm{in}}},\varepsilon_{\mathrm{in}}\rangle$-IV-fidelitous if
@@
Given $F$, $x_{s}\!=\!G_{\Theta}(SD)$ is $\langle F,\operatorname{D_{\mathrm{out}}},\varepsilon_{\mathrm{out}}\rangle$-OV-fidelitous if
@@
Beyond input/output matching, one can assess whether the SUT processes synthetic inputs through similar internal mechanisms. Given a deep neural network (DNN) $F$ with intermediate layers indexed by $\ell$, and using a distance function $\operatorname{D}_{\mathrm{lf}}$ and tolerance $\varepsilon_{\mathrm{lf}}$:
@@
IV Decisive-Feature Fidelity (DFF)
Our framework, illustrated in Fig. 1, consists of three main components: (1) Metric Definition: We formalize DFF to quantify mechanism parity alongside IV and OV fidelity. (2) Estimation: An XAI explainer extracts decisive features from both real and synthetic inputs to compute the DFF metric. (3) Calibration: A feedback loop uses the DFF metric to optimize the generator’s parameters $\Theta$, minimizing the mechanism gap. The following subsections detail each component.
@@ IV-A DFF Metric Definition
We instantiate decisive features via an XAI explainer $\operatorname{H}$ and compare them using a distance function $\operatorname{D}$. Note, we exclude LF fidelity ($\varepsilon_{\text{lf}}$) from the top-level acceptance criteria as DFF provides a more specific measure of internal mechanism alignment relevant to the task. The thresholds $\varepsilon_{\text{in}}$, $\varepsilon_{\text{out}}$, and $\varepsilon_{\text{dff}}$ are user-defined hyperparameters. In practice, they are derived empirically from a calibration set (e.g., using the 95th percentile of distances between matched real pairs) or set based on industry standards (e.g., OV error $5\%$). Formally, $x_{s}$ is $\langle F,\operatorname{H},\operatorname{D},\varepsilon_{\text{dff}}\rangle$-DFF-fidelitous with respect to $x_{r}$ if
@@
$\text{Pass-Rate}=\frac{1}{N}\sum_{i=1}^{N}\mathbb{I}\Bigl(\operatorname{D}\bigl(\operatorname{H}(F(x_{s,i})),\operatorname{H}(F(x_{r,i}))\bigr)\leq\varepsilon_{\text{dff}}\Bigr).$ (4)
$\mathbb{I}(x)=\begin{cases}1,&\text{if }x=\mathrm{TRUE},\\ 0,&\text{otherwise}.\end{cases}$
@@
IV matches pixels, OV matches outputs, LF matches broad internals; DFF matches decisive evidence and can diverge under mechanism shifts (right output for wrong reason).